OpenStudy (anonymous):

using the number 4 four times, come up with the answer 13 using addition, subtraction, multiplication, or square roots.

OpenStudy (anonymous):

no division?

OpenStudy (anonymous):

yes, you can use division...sorry

OpenStudy (anonymous):

only use 4 four times?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can't use any other numbers

OpenStudy (anonymous):

can you use the #44?

OpenStudy (anonymous):

no...only 4 fours.

OpenStudy (anonymous):

the only way i can think to do it is .. \[\frac{4+4!+4!}{4}\]

OpenStudy (anonymous):

Sorry, what does the ! stand for?

OpenStudy (anonymous):

factorial .. 4! = 4*3*2*1 = 24

OpenStudy (anonymous):

no, I doubt that's it. my son's only a seventh grader and hasn't used factorials yet.

OpenStudy (anonymous):

i can only do it that way or using 44

OpenStudy (anonymous):

okay, well, let's try the number 44. How do you go about that?

OpenStudy (anonymous):

\[\frac{44}{4} + \sqrt{4}\]

OpenStudy (anonymous):

thanks sweetie, how about trying the same thing with the number 11 now?

OpenStudy (anonymous):

\[\frac{44}{\sqrt{4}+\sqrt{4}}\]

OpenStudy (anonymous):

oh, you are so great! Thank you, dear. You just helped us out a great deal.

OpenStudy (anonymous):

anytime.